a chemist needs to mix up 50 gallons of a 60% acid solution. she only has 50% and 75% solutions on hand. how many gallons should she mix together?
Answer #1

Solution:-

Given that, a chemist has only 50% and 75% solutions on hand. She needs to mix up 50 gallons of a 60% acid solution.

We have to determine the number of gallons she should mix together inorder to obtain 50 gallons of a 60% acid solution.

Let,

T = gallons of the 50% solution

y = gallons of the 75% solution

The total volume of the new mixture is 50 gallons.

Therefore,

x+y=50\; \; ....equation(1)

According to question,

Each gallon of 50% solution contributes 0.5 gallon of acid, and each gallon of the 75% solution contributes 0.75 gallon of acid. The final mixture needs to have a concentration of 60% acid.

Hence,

0.5.0 + 0.75y = 0.6 x 50

0.5x+0.75y = 30\; \; ....equation(2)

From\; \; equation(1):

x+y=50

x = 50-y

Put\; \; x= 50-y\; \; in\; \; equation(2)

0.5x+0.75y = 30

0.5(50-y)+0.75y = 30

0.5\times (50)-0.5\times (y)+0.75y = 30

25-0.5y+0.75y = 30

25+0.25y = 30

0.25y = 30-25

0.25y = 5

y = \frac{5}{0.25}

\boldsymbol{y = 20}\; \; gallons

Put\; \; y= 20\; \; in\; \; equation(1)

x+y = 50

x+20 = 50

x= 50-20

\boldsymbol{x= 30}\; \; gallons

Hence,

She should mix \boldsymbol{30} gallons of the 50% solution and \boldsymbol{20} gallons of the 75% solution.